Caluculate the volume at NTP occupied by 6.25g of nitrogen

Grade:11

2 Answers

Vikas TU
7 years ago
First Calculate the number of moles of Nitrogen,
n = 6.25/28 
At NTP,
V = n*22.4
put the n value and get the volume in litre.
kishan
6 years ago
no. of mole of 6.25g nitrogen =6.25/28 [equation 1]at ntp conditionno. of mole=V/22.4 [equation 2]by equation 1 and 2V/22.4=6.25/28V=6.25/28×22.4V=5ans V=5lit